### v1

# Enhancing Answer Reliability Through Inter-Model Consensus of Large Language Models

Abstract
We explore the collaborative dynamics of an innovative language model interaction system involving advanced models such as GPT-4-0125-preview, Meta-LLaMA-3-70B-Instruct, Claude-3-Opus, and Gemini-1.5-Flash. These models generate and answer complex, PhD-level statistical questions without exact ground-truth answers. Our study investigates how inter-model consensus enhances the reliability and precision of responses. By employing statistical methods such as chi-square tests, Fleiss’ Kappa, and confidence interval analysis, we evaluate consensus rates and inter-rater agreement to quantify the reliability of collaborative outputs. Key results reveal that Claude and GPT-4 exhibit the highest reliability and consistency, as evidenced by their narrower confidence intervals and higher alignment with question-generating models. Conversely, Gemini and LLaMA show more significant variability in their consensus rates, as reflected in wider confidence intervals and lower reliability percentages. These findings demonstrate that collaborative interactions among large language models (LLMs) significantly improve response reliability, offering novel insights into autonomous, cooperative reasoning and validation in AI systems.
Keywords: Large Language Models, Collaborative Intelligence, Answer Validation, Game Theory, Statistical Analysis
1 Introduction
The emergence of large language models (LLMs) has transformed natural language processing and artificial intelligence, enabling machines to perform complex language tasks with unprecedented proficiency. Models such as OpenAI’s GPT-4, Meta’s LLaMA series, Anthropic’s Claude, and Google’s Gemini have demonstrated remarkable capabilities in generating human-like text, understanding context, and even exhibiting reasoning abilities Mann et al. (2020); Touvron et al. (2023). These advancements have opened new avenues for automated knowledge generation and validation in specialized domains Kojima et al. (2022).
However, a significant challenge remains in validating the outputs of these models when predefined correct answers or ground truths are unavailable. This issue is particularly critical in specialized fields like advanced statistics, where the complexity and nuance of questions cause manual verification to be resource-intensive and impractical Jiao et al. (2023). Traditional validation methods that rely on comparing model outputs against predetermined correct answers become insufficient in such contexts Hendrycks et al. (2020).
1.1 Research Context and Motivation
Leveraging multiple LLMs to enhance answer reliability through collaboration and consensus presents a promising approach. By harnessing the collective intelligence of several models, it becomes possible to approximate correctness and identify consensus even in the absence of ground truth He et al. (2022). This methodology draws inspiration from ensemble learning techniques in machine learning Dietterich (2000) and the wisdom of crowds phenomenon in human collective intelligence Mennis (2006).
Recent studies indicate that different LLMs may exhibit complementary strengths and weaknesses in their reasoning capabilities Taylor et al. (2022). This observation suggests that a collaborative framework leveraging multiple models could overcome individual limitations and yield more reliable answers. However, the dynamics of such inter-model collaboration, especially in specialized domains like statistics, remain largely unexplored.
1.2 Theoretical Framework
This research synthesizes three fundamental theoretical frameworks: collective intelligence theory, distributed cognition framework, and consensus formation models to understand and analyze the collaborative dynamics among LLMs in statistical reasoning tasks.
Collective Intelligence Theory Lévy (1997) provides a foundational basis for understanding how multiple agents can collaborate to achieve superior outcomes compared to individual performance. In the context of LLMs, this theory suggests that different LLM architectures (e.g., GPT-4, Claude-3, LLaMA, Gemini) offer varied approaches to problem-solving, potentially leading to more robust solutions Hong and Page (2004). It informs our approach to combining individual model outputs through structured consensus formation Woolley et al. (2010). Additionally, each model’s independent processing of questions helps maintain solution diversity and reduces cascading errors Vercammen et al. (2019).
The Distributed Cognition Framework Hutchins (1995) offers theoretical grounding for understanding how cognitive processes can be distributed across multiple artificial agents. This framework is particularly relevant because different models may encode complementary aspects of statistical knowledge, leading to more comprehensive problem-solving capabilities Du et al. (2024). Complex statistical reasoning tasks can be decomposed and processed across multiple models, potentially improving overall solution quality Liu et al. (2024). Furthermore, the framework helps explain how information flows between models during collaborative problem-solving Naik (2024).
Mathematical models of consensus formation provide the theoretical basis for understanding how agreement emerges among multiple decision-making agents Friedkin (1990). These models inform the fundamental dynamics of collaborative AI systems through three key aspects: the dynamics of agreement, which describe how consensus evolves; weighted influence mechanisms, which define the relative impact of each agent in the consensus process Olfati-Saber et al. (2007); and the conditions for convergence that ensure the stability of collaborative solutions Baronchelli (2018). Recent work by Bahrami et al. (2010) has demonstrated how these theoretical frameworks can be effectively applied to multi-agent decision-making systems, offering a robust foundation for understanding collaborative behavior in artificial intelligence.
Integrating collective intelligence, distributed cognition, and consensus formation frameworks provides a comprehensive foundation for analyzing collaborative AI systems Hutchins (1995); Woolley et al. (2010). This integration can be formally expressed as:
$R=f(\text{CI},\text{DC},\text{CF})$ (1)
where $R$ represents the reliability of collaborative outcomes, integrated across Collective Intelligence (CI), Distributed Cognition (DC), and Consensus Formation (CF) principles Patrikalakis et al. (1999). Through this unified approach, we can better understand how multiple AI models collaborate and generate reliable solutions in complex reasoning tasks.
Our analytical framework synthesizes quantitative and qualitative methods derived from these theoretical perspectives Klein et al. (2006). The quantitative aspects draw from consensus formation models, providing metrics for measuring agreement and convergence in multi-agent systems Mennis (2006).
The validation methodology emerging from this integrated framework emphasizes the importance of multiple perspectives in evaluating model outputs Kittur and Kraut (2008). By considering various theoretical viewpoints, we can better assess the reliability and robustness of collaborative AI solutions Malone and Bernstein (2022). This comprehensive theoretical foundation grounds our investigation in established principles and provides new insights into the unique challenges posed by AI collaboration in statistical reasoning tasks Page (2008).
1.3 Research Gap
While existing research has extensively examined individual LLM performance Ahn et al. (2024) and basic ensemble methods Huang et al. (2024), there is a notable gap in understanding how multiple state-of-the-art LLMs can collaboratively validate complex knowledge in the absence of ground truth. This gap is particularly significant in specialized academic domains where:
  • The complexity of questions requires sophisticated reasoning capabilities.
  • Manual validation by human experts is time-consuming and costly.
  • Traditional automated validation methods are insufficient.
  • The dynamic nature of knowledge makes maintaining up-to-date ground truths challenging.
Addressing these challenges necessitates novel approaches that leverage collaborative validation among LLMs, providing a pathway toward reliable, scalable, and efficient knowledge validation.
1.4 Objectives of the Research
This study aims to explore and understand the collaborative dynamics among different LLMs in statistical reasoning tasks. The primary objectives are to examine how answers from different LLMs align or differ when responding to complex statistical questions generated by one of the models. This includes investigating how different LLMs complement each other in collaborative statistical reasoning and understanding the patterns of their interactions. Then, develop a framework for evaluating the reliability of answers based on model consensus, especially in the absence of ground truth. This involves quantifying the reliability of collaborative outcomes without predefined correct answers and identifying metrics that best capture the quality of inter-model agreement. Also, statistical techniques such as chi-square tests and Fleiss’ Kappa are used to assess the significance of consensus rates and measure inter-model agreement. This objective supports validating collaborative mechanisms and the effectiveness of consensus formation among multiple LLMs.
1.5 Significance and Contributions
This research makes several significant contributions to the field:
  1. 1.
    Introduces a novel framework for validating complex knowledge through multi-model collaboration.
  2. 2.
    Provides empirical evidence for the effectiveness of collaborative validation approaches.
  3. 3.
    Establishes benchmarks for future research in LLM-based knowledge validation.
The findings of this study have implications in several areas, such as educational technology and automated assessment systems, research validation in specialized academic domains, development of more reliable AI-powered knowledge systems, and improving our understanding of how artificial systems can collaborate and exhibit collective intelligence to solve complex problems to name a few.
2 Literature Review
AI trends have moved from advances in image processing and computer vision Voulodimos et al. (2018); Azad et al. (2024) to progress in reinforcement learning Wiering and Van Otterlo (2012); Jebellat et al. (2021, 2024) and natural language processing Chowdhary and Chowdhary (2020); Pillai (2023), leading to today’s focus on LLMs and generative AI.
LLMs have achieved significant milestones in natural language understanding and generation Radford et al. (2019). The GPT series, notably, has demonstrated remarkable capabilities in few-shot and zero-shot learning scenarios Achiam et al. (2023). Recent advancements in model architectures and training approaches have led to increasingly sophisticated systems, such as Claude-3 Anthropic (2024) and Gemini Team et al. (2023), which exhibit enhanced reasoning abilities in specialized domains.
The concept of collaborative intelligence among AI agents has emerged as a promising approach for enhancing problem-solving capabilities Dafoe et al. (2020). Recent studies have shown that model collaboration can significantly improve reasoning capabilities through cross-model validation Yin et al. (2023). Additionally, it enhances robustness, particularly in tackling complex problem-solving tasks Sun et al. (2023); Jebellat and Jebellat (2024); Gholami Davoodi et al. (2024), and leads to more reliable outputs by leveraging consensus mechanisms.
Integrating collaborative strategies in artificial intelligence (AI) systems has consistently demonstrated enhanced performance across various domains. For instance, Lu et al. (2024) conducted a comprehensive survey highlighting that merging, ensembling, and cooperative approaches among LLMs lead to superior outcomes in natural language processing tasks. In the medical field, Rezk and Selim (2024) reviewed metaheuristic-based ensemble learning methods, emphasizing their effectiveness in improving diagnostic accuracy and treatment planning. Similarly, in mathematical problem-solving, researchers at MIT developed a multi-AI collaboration framework that enhances reasoning and factual accuracy in LLMs, resulting in more reliable solutions to complex mathematical queries MIT News (2023). In software engineering, ensemble methods have been shown to improve code defect detection and software quality assurance processes Gupta et al. (2022). These findings collectively highlight the significant advantages of collaborative AI approaches, particularly as LLMs continue to evolve and scale.
The collaboration of AI models raises critical ethical questions regarding transparency, accountability, and bias propagation Bender et al. (2021). Ensuring that collaborative AI systems operate ethically is crucial, especially in the absence of ground-truth verification Mittelstadt (2019).
A primary concern is the potential for bias amplification in collaborative systems. Raghavan et al. (2020) demonstrated that model ensembles can compound existing biases, with their study showing significant increases in gender and demographic biases when multiple models interact. Building on this work, Hashimoto et al. (2018) established frameworks for measuring and mitigating such cumulative biases in machine learning systems.
Transparency presents another significant challenge in collaborative AI systems. Doshi-Velez et al. (2017) identified critical areas, including decision attribution complexity, interpretability of consensus mechanisms, and accountability frameworks. These challenges become particularly essential in high-stakes applications, where understanding model decisions is crucial Rudin (2019).
Recent case studies have highlighted these ethical considerations across various domains. In healthcare applications, Larrazabal et al. (2020) examined ethical implications in medical imaging diagnosis, finding that collaborative systems required additional safeguards for patient privacy and decision transparency. Similarly, in educational contexts, Holstein et al. (2019) investigated fairness in automated assessment systems, revealing the need to carefully calibrate collaborative systems to ensure equitable evaluation across diverse student demographics. Gebru et al. (2021) further emphasized the importance of comprehensive documentation and transparency in AI systems, mainly when multiple models work in conjunction.
3 Methodology
3.1 Research Design
This study employs a mixed-methods approach to investigate the collaborative dynamics among LLMs in statistical reasoning tasks. The experimental design incorporates quantitative analysis of model consensus patterns. The study framework can be formalized as follows:
$S=\{M,Q,A,V\},$ (2)
where $M$ represents the set of models, $Q$ the question generation process, $A$ the answering process, and $V$ the validation mechanisms.
3.2 Overview
We examine how multiple LLMs collaborate to generate and validate complex, PhD-level statistical multiple-choice questions (MCQs) without ground-truth answers. A total of $N=100$ MCQs were generated and answered. One LLM acted as the question generator for each question, while the remaining three independently provided answers and justifications. Question generator and answerer roles were rotated among the models to ensure interchangeability and mitigate model-specific biases.
3.3 Language Models Used
We utilized four state-of-the-art LLMs, each with distinct architectural characteristics:
Table 1: Descriptions of the Four Language Models Used
Model Description
GPT-4-0125-preview An advanced model renowned for its reasoning capabilities and sophisticated natural language understanding OpenAI (2023).
Meta-LLaMA-3-70B-Instruct A 70-billion-parameter model optimized for instruction following, featuring enhanced few-shot learning capabilities Touvron et al. (2023).
Claude-3-Opus Designed to provide safe and helpful responses focusing on complex tasks Anthropic (2024).
Gemini-1.5-Flash A multimodal model integrating language and vision Team et al. (2023).
Each model was accessed via its respective API, using default hyperparameters unless specified otherwise. All models operated under identical conditions to maintain consistency.
3.4 Question Generation and Answering Processes
To ensure the creation of challenging and diverse multiple-choice questions (MCQs) suitable for PhD-level statistics, we designed an integrated framework combining question generation and independent answering LLMs. This framework also incorporated strategies to mitigate potential biases in the question generation and answering processes.
The question generation phase leveraged a comprehensive concept map of advanced statistical topics, as illustrated in table 2. This table encompasses a diverse set of topics $\mathcal{T}$ and subtopics $\mathcal{S}$. For each question, a topic $t_{i}\in\mathcal{T}$ and subtopic $s_{i}\in\mathcal{S}$ were randomly selected to ensure uniform coverage and diversity. The question-generating LLM received a carefully designed prompt $P_{q}$:
”Generate a challenging PhD-level multiple-choice question in the field of [Topic], focusing on [Specific Concept]. The question should have four answer options labeled A, B, C, and D, with only one correct answer. Ensure the question tests deep understanding and critical thinking skills.”
Here, the placeholders [Topic] and [Specific Concept] were populated based on the selected topic $t_{i}$ and subtopic $s_{i}$ from the table 2. The generated output included the question $Q_{i}$, the answer set $\{A_{i},B_{i},C_{i},D_{i}\}$, the correct answer $A_{i}^{c}$, and an explanation $E_{i}$. Notably, $A_{i}^{c}$ and $E_{i}$ were retained for analysis but withheld from the answering models to replicate real-world conditions where ground-truth answers are unavailable.
To mitigate potential biases, we implemented several strategies. First, topic diversity was ensured by including various topics and subtopics, preventing overrepresenting specific areas. Second, the crafted prompts were neutral and avoided introducing leading language or biases, ensuring no influence on the models’ responses in a particular direction. Finally, all generated content was manually reviewed to identify and exclude inappropriate or biased material, thereby preserving the integrity and fairness of the dataset.
Table 2: Comprehensive overview of statistical topics and their key subtopics. This served as the foundation for question topic selection.
Topic Subtopics
Bayesian inference Prior distributions, Posterior updating, Bayes factors, Conjugate priors
Markov Chain Monte Carlo methods Metropolis-Hastings algorithm, Gibbs sampling, Convergence diagnostics, Sampling efficiency
Time series analysis ARIMA models, Stationarity, Seasonality, Spectral analysis
Multivariate statistics Principal Component Analysis, Factor analysis, Canonical correlations, Multivariate normal distribution
Hypothesis testing Type I and Type II errors, Power analysis, Non-parametric tests, Multiple comparisons correction
Non-parametric methods Kernel density estimation, Bootstrap methods, Spline regression
Survival analysis Cox proportional hazards model, Kaplan-Meier estimator, Censoring, Hazard functions
Experimental design Randomization techniques, Blocking and confounding, Factorial designs, Response surface methodology
Regression analysis Generalized linear models, Heteroscedasticity, Collinearity, Model selection criteria
Statistical learning theory Overfitting and underfitting, Regularization techniques, Bias-variance tradeoff, Cross-validation
In the answering phase, three distinct LLMs independently attempted the generated questions. Each answering model received the question $Q_{i}$ and answer set $\{A_{i},B_{i},C_{i},D_{i}\}$ alongside the following prompt $P_{a}$:
”Please read the following Ph.D.-level statistics question and select the most appropriate answer (A, B, C, or D). Provide a detailed justification for your selection, explaining your reasoning and any relevant statistical principles.”
Each model independently selected an answer $a_{ij}$ for question $Q_{i}$, where $j\in\{1,2,3\}$ indexes the answering models, and provided a justification $J_{ij}$. The models were isolated from one another, ensuring no exchange of information about answers or justifications to prevent bias or collusion. This isolation, combined with the neutral prompts and diverse question topics, further minimized bias in the answering process.
3.5 Inter-Model Consistency Analysis
We performed an inter-model consistency analysis to assess the agreement among the answering LLMs and evaluate the reliability of their consensus. For each question $Q_{i}$, we collected and organized data into a comprehensive dataset $D_{i}$, defined as:
$D_{i}=\{Q_{i},\{A_{i},B_{i},C_{i},D_{i}\},\{a_{i1},a_{i2},a_{i3}\},\{J_{i1},J_{i2},J_{i3}\}\}$ (3)
This dataset encompassed the generated questions $Q_{i}$, their corresponding multiple-choice options $\{A_{i},B_{i},C_{i},D_{i}\}$, the responses from each participating LLM—including selected answers $a_{ij}$ and justifications $J_{ij}$—and metadata such as model identifiers, timestamps, and relevant model-specific parameters. This structured approach facilitated quantitative and qualitative analysis of the reasoning processes and agreement patterns among models Bommasani et al. (2021).
We analyzed the alignment and divergence in the selected answers $a_{ij}$ to quantify inter-model consistency. The degree of consensus was categorized as follows:
  • Full Agreement: All three models selected the same answer.
  • Partial Agreement: Two models selected the same answer, while one differed.
  • No Agreement: All models selected different answers.
This classification provided a robust framework for evaluating agreement levels and understanding the variability in model decision-making, offering insights into their collective and individual reasoning patterns.
3.6 Answer Validation Mechanisms
Given the absence of ground-truth answers, we employed several mechanisms to validate the models’ answers based on inter-model consensus.
Majority Vote, Reliability, and Confidence Interval
To determine the consensus answer $A_{i}^{\text{cons}}$ for a given question $Q_{i}$, we utilized a majority voting mechanism over the responses from multiple LLMs. Formally, the consensus answer is defined as:
$A_{i}^{\text{cons}}=\arg\max_{k\in\{A,B,C,D\}}\sum_{j=1}^{3}\delta(a_{ij},k),$ (4)
where $a_{ij}$ denotes the response provided by the $j$-th LLM for question $Q_{i}$, and $\delta(a_{ij},k)$ is the Kronecker delta function:
$\delta(a_{ij},k)=\begin{cases}1,&\text{if }a_{ij}=k,\\ 0,&\text{otherwise.}\end{cases}$
This formulation assigns a frequency-based score to each answer option $k$, with the majority-selected response designated as the consensus answer. Both partial and complete agreements are considered valid, forming the basis for subsequent analyses.
In the absence of ground truth labels, we introduced a reliability metric to assess the trustworthiness of the consensus answers. Reliability measures the alignment between the consensus response and the answer provided by the LLM that generated the question. Let $A_{i}^{\text{LLM-q}}$ denote the querying LLM’s response for question $Q_{i}$. The reliability score $R_{i}$ is computed as:
$R_{i}=\begin{cases}1,&\text{if }A_{i}^{\text{cons}}=A_{i}^{\text{LLM-q}},\\ 0,&\text{otherwise.}\end{cases}$ (5)
A higher reliability score indicates a more substantial alignment between the consensus answer and the querying LLM’s response, increasing confidence in the consensus’s validity. This measure is precious in settings where ground truth data is unavailable.
To evaluate the robustness of consensus rates across LLMs, we calculated confidence intervals (CIs) using a bootstrap resampling approach. Confidence intervals provide a probabilistic range for the actual consensus rate, capturing variability in the observed data. Narrow CIs indicate high reliability, while wider CIs reflect more significant uncertainty.
Bootstrap resampling enables CI estimation without assumptions about the underlying data distribution. Specifically, $B$ bootstrap samples are generated by sampling the original dataset with replacement. The mean agreement rate is computed for each sample, constructing a bootstrap distribution of the consensus rate. The $\alpha/2$-th and $21-\alpha/$ 2-th percentiles of this distribution define the lower and upper CI bounds, respectively. For a 95% CI, these correspond to the 2.5th and 97.5th percentiles.
By incorporating confidence intervals, we mitigate uncertainty in agreement rates and establish a statistical basis for comparing LLMs. Overlapping CIs suggest no significant differences between models, whereas non-overlapping intervals indicate statistically significant differences. This methodology ensures a robust and interpretable evaluation of model reliability, mainly when ground truth labels are unavailable.
To further validate reliability and assess statistical significance, we employed the following additional statistical measures:
Chi-Square Test of Independence
A chi-square test was used to evaluate whether the distribution of selected answers deviated significantly from random chance. The test statistic was calculated as:
$\chi^{2}=\sum_{k=1}^{K}\frac{(O_{k}-E_{k})^{2}}{E_{k}},$ (6)
Where:
  • $O_{k}$ is the observed frequency of answer choice $k$ across all models and questions.
  • $E_{k}$ is the expected frequency of answer choice $k$, assuming uniform random selection, $E_{k}=\frac{N\times n_{j}}{K}$.
  • $K=4$ is the number of answer choices.
  • $N$ is the total number of questions.
  • $n_{j}=3$ is the number of answering models.
To determine statistical significance, the computed $\chi^{2}$ value was compared against the chi-square distribution with $K-1$ degrees of freedom.
Fleiss’ Kappa Coefficient
Fleiss’ kappa $\kappa$ was employed to quantify the agreement among the models beyond chance. It is defined as:
$\kappa=\frac{\overline{P}-\overline{P_{e}}}{1-\overline{P_{e}}},$ (7)
Where:
  • $\overline{P}$ is the mean observed agreement among the models.
  • $\overline{P_{e}}$ is the mean agreement expected by chance.
The value of $\kappa$ ranges from $-1$ (perfect disagreement) to $1$ (perfect agreement), with 00 indicating no agreement beyond chance.
These mechanisms collectively ensured a robust framework for validating model responses and assessing the reliability of the generated consensus.
4 Results
4.1 Inter-Model Consistency Analysis
We assessed the level of agreement among the models—GPT-4, LLaMA, Gemini, and Claude—for each generated question. In each experiment, one model generated $100$ questions, and the remaining three independently provided answers without collaboration. The implantation and data are available here Amiri-Margavi (2024).
Figure 1: Consensus rate overview showing the levels of agreement among the models when different LLMs generate the questions. The x-axis represents the question-generating LLM model, and the y-axis indicates the percentage of responses in each agreement category.
[…]
These findings suggest that inter-model consistency depends significantly on the question-generating model. The higher agreement levels observed with Claude and GPT-4 may be attributed to their advanced training and more sophisticated understanding of complex statistical concepts. Conversely, the lower consensus rates with LLaMA and Gemini could be influenced by differences in their model architectures, training data, or interpretation nuances. This variation underscores the impact of the question generator on the collaborative performance of LLMs in specialized domains.
To evaluate the reliability and precision of answers generated by LLM) in scenarios lacking ground-truth data, we employed two complementary mechanisms: majority voting and confidence interval analysis. These approaches quantify inter-model agreement and assess alignment with a reference answer implicitly provided by the question-generating LLM.
Using the majority voting approach, an answer is considered reliable if at least two other LLMs agree with the question-generating LLM’s response. Figure 2 presents the majority vote and reliability percentages for various LLMs. Key observations reveal that majority vote percentages are consistently higher than reliability percentages across all models, indicating strong inter-model agreement but fewer instances where the models’ answers align with the reference answer from the question-generating LLM.
Figure 2: Majority vote and reliability percentages across different models. The majority vote percentage indicates instances where at least two models agree, while the reliability percentage reflects instances where two or more models agree with the question-generating LLM’s answer.
[…]
This analysis highlights the utility of majority voting in quantifying inter-model agreement and reliability. The gap between majority vote and reliability percentages highlights the challenges of aligning responses with a reference answer, particularly in the absence of ground truth data. These findings demonstrate the effectiveness of majority voting as a plausible mechanism for validating answers and assessing model performance when definitive correctness cannot be directly measured.
To further evaluate the precision and reliability of inter-model agreement, we calculated confidence intervals (CIs) for the consensus rates of each model. Confidence intervals provide a statistical range within which the consensus rate is likely to fall, enabling an assessment of variability and robustness. Table 3 summarizes each model’s lower and upper bounds of the CIs.
Table 3: Confidence intervals for consensus rates across models.
Model Lower Bound Upper Bound
Gemini 0.60 0.78
Claude 0.80 0.93
GPT-4 0.75 0.90
LLaMA 0.55 0.74
Analysis of the confidence intervals reveals distinct patterns in the precision and reliability of the models. Claude demonstrates the narrowest CI (0.80–0.93), suggesting high reliability and minimal variability in its consensus rates. GPT-4 follows closely with a CI of (0.75–0.90), indicating similarly high reliability but slightly increased variability. In contrast, Gemini and LLaMA exhibit broader CIs, spanning (0.60–0.78) and (0.55–0.74), respectively, reflecting more significant variability and reduced precision in their consensus rates.
These findings provide quantitative evidence of inter-model performance differences. Models with narrower CIs, such as Claude and GPT-4, offer more reliable consensus rates, making them more suitable for high-precision tasks. Conversely, the wider CIs observed for Gemini and LLaMA highlight potential inconsistencies and need further refinement. Confidence interval analysis thus offers a robust framework for assessing model agreement and reliability without ground truth data.
4.2 Statistical Significance Testing
We employed two critical approaches to assess the statistical significance of agreement among the models: chi-square testing and Fleiss’ Kappa analysis. These methods quantified the extent to which the models’ responses deviated from random chance and evaluated their inter-rater agreement.
The chi-square test was performed to determine whether the observed agreement among the models occurred by chance. The null hypothesis assumed that the distribution of answers was random, indicating no meaningful consensus. The test statistic was calculated based on eq: 6, in which the number of choices $K=4$ and $n_{j}=3$ is the number of answering models. The p-values obtained from the test are summarized in Table 4.
Table 4: Chi-square test p-values indicating the statistical significance of agreement for each model.
Model Gemini Claude GPT-4 LLaMA
p-value $4.29\times 10^{-26}$ $4.29\times 10^{-46}$ $0.00546$ $1.67\times 10^{-10}$
All p-values were below the standard significance level ($\alpha=0.01$), leading us to reject the null hypothesis. The extremely small p-values for Claude, Gemini, and LLaMA indicate statistically solid significance, confirming that their agreements are highly unlikely to be due to random chance. GPT-4 also shows statistical significance, though its p-value is less extreme than the others. These results support the hypothesis that the models’ agreements reflect meaningful consensus rather than random behavior.
Fleiss’Kappaa was further employed to evaluate inter-rater agreement among the models. This metric provides a robust measure of consistency in responses. Table 5 presents the kappa values and their interpretations.
Table 5: Fleiss’ Kappa values indicating the level of agreement among models.
Model Kappa Value Interpretation
Gemini 0.2811 Fair agreement
Claude 0.7160 Substantial agreement
GPT-4 0.4275 Moderate agreement
LLaMA 0.5572 Moderate agreement
The kappa values reveal varying levels of agreement across the models. Claude demonstrated substantial agreement, reflecting strong consistency with other models. GPT-4 and LLaMA Combiningderate agreement, indicating reasonable consistency but some response variability. Conversely, Gemini showed fair agreement, suggesting noticeable differences in its responses compared to the other models.
These results highlight potential differences in model architectures, training data, or reasoning approaches. For instance, Gemini’s lower agreement could stem from less accurate or diverse data sources, while Claude’s substantial agreement may be attributed to robust training on well-aligned datasets. By combining chi-square testing and Fleiss’ Kappa analysis, we comprehensively evaluate the statistical significance and inter-model agreement, demonstrating the reliability and coherence of LLM-generated answers.
5 Discussion
5.1 Enhancing Answer Reliability Through Collaboration
The high consensus rate among the models, as demonstrated in both the majority vote and reliability analyses, underscores the potential of collaborative dynamics in enhancing answer reliability. The substantial Fleiss’ Kappa values confirm that the agreement among the models is statistically significant and not due to chance. This collective agreement suggests that the models possess a shared understanding of statistical concepts, even without ground-truth answers.
In scenarios where expert validation is unavailable—such as rapidly evolving domains or complex problem-solving contexts—leveraging multiple LLMs for consensus-based validation provides a valuable proxy for correctness. For instance, questions where Claude and GPT-4 exhibit high agreement can be considered more reliable, whereas questions with lower agreement or broader confidence intervals (e.g., those involving LLaMA) may require further assessment. The p-values from the chi-square test further support this analysis, allowing researchers to identify questions where models show strong alignment versus those with more significant variability. This insight can guide future efforts to refine questions with high disagreement rates, which may be inherently ambiguous or beyond the current capabilities of LLMs.
5.2 Implications for AI and Education
Our findings significantly impact applying LLMs in AI-driven education and content generation. In collaborative learning, LLMs can enhance the quality of educational content by validating and refining answers collaboratively, even without human-provided ground-truth answers. This capability could revolutionize automated learning systems, enabling them to independently generate and validate high-quality content. Such systems could provide immediate, reliable feedback to learners, adapt to new information swiftly, and support educators by handling routine assessment tasks.
5.3 Limitations
This study acknowledges several limitations that may impact the findings and interpretations. One significant limitation is the potential ”model similarities” or ”model homogeneity.” Despite differences in architecture and training, the analyzed LLMs may share overlapping training data, leading to correlated errors and overestimating consensus. This overlap could result in models reinforcing each other’s misconceptions rather than providing independent validation.
A second limitation is the ”lack of a human benchmark.” Without expert human evaluation, it remains challenging to assess the absolute correctness of the consensus answers definitively. The reliability metrics employed, while meaningful, serve as subjective proxies and cannot fully substitute for expert judgment.
Third, the study highlights the risk of ”bias propagation.” If all models share similar misconceptions, they may propagate and reinforce incorrect answers through agreement, which could be particularly problematic in specialized or sensitive domains where accuracy is critical.
Finally, this study is constrained by its ”static evaluation” nature. It captures a snapshot of the models’ capabilities at a single point in time and does not account for potential updates, improvements, or degradations in model performance over time. Future work should consider dynamic evaluation to track evolving model capabilities better.
5.4 Conclusion
[…]
These findings prove that collaborative interactions among large language models (LLMs) can improve response reliability through consensus mechanisms. The study highlights the potential of these methodologies for applications in AI-driven education, automated validation systems, and collaborative AI frameworks. For example, leveraging highly reliable models such as Claude and GPT-4 in cooperative settings could enhance the accuracy and trustworthiness of AI-generated responses in high-stakes domains.
Future research should explore integrating models with diverse architectures and training datasets to mitigate the risk of correlated errors and investigate the impact of model heterogeneity on consensus rates. Further analysis of bias propagation is also critical to ensure collaborative systems do not inadvertently reinforce shared misconceptions among models. Addressing these directions will significantly improve consensus-based validation frameworks’ robustness, scalability, and reliability, laying the groundwork for more effective and trustworthy AI systems.